What is the volume V of a sample of 2.90 mol of copper? The atomic mass of copper (Cu) is 63.5 g/mol, and the density of copper is 8.92×103kg/m3. Express your answer in cubic centimeters to 3 significant figures.
Answer #1

Here ,

n = 2.90 mol

total mass of 2.90 moles , m = 2.90 * 63.5 gm

m = 184.15 gm = 0.185 Kg

volume of sample = 0.185/(8.92 *10^3)

volume of sample = 2.074 *10^-5 m^3

the volume of the sample is 2.074 *10-5 m^3